How many ounces of a 5% alcohol solution must be mixed with 15 ounces of a 10% alcohol solution to make a 6% alcohol solution
Answer by stanbon(75887)

How many ounces of a 5% alcohol solution must be mixed with 15 ounces of a 10% alcohol solution to make a 6% alcohol solution
----
Equation:
alch + alch = alch
0.10*15 + 0.05x = 0.06(15+x)
-------
10*15 + 5x = 6*15 + 6x
x = 4*15 = 60 ounces (amt of 5% solution needed)
-------
Cheers,
Stan H.
-----------